fraction numerator x squared plus x plus 1 over denominator x minus 1 end fraction



1
Expert's answer
2021-10-11T01:42:10-0400
"\dfrac{x^2+x+1}{x-1}=\dfrac{x^2-x+2x-2+3}{x-1}"

"=x+2+\dfrac{3}{x-1}, x\not=1"

"\dfrac{x^2+x+1}{x-1}=x+2+\dfrac{3}{x-1}, x\not=1"



"\int\dfrac{x^2+x+1}{x-1}dx=\int(x+2+\dfrac{3}{x-1})dx"

"=\dfrac{x^2}{2}+2x+3\ln(|x-1|)+C"

"\int\dfrac{x^2+x+1}{x-1}dx=\dfrac{x^2}{2}+2x+3\ln(|x-1|)+C"